OpenStudy (anonymous):

A merry-go-round travels in a circle at a velocity of 3.0 m/s. The diameter of the merry-go-round is 3.0 meters. What is the centripetal acceleration of the merry-go-round? I thought this was a simple problem but my answer isn't matching up with any of the options so I was hoping to see if someone else got something other than me (to see if it was my error or the book's)

OpenStudy (aaronq):

did you do: \(a_r=\dfrac{v^2}{r}=\dfrac{(3~m/s)^2}{1.5 ~m}=\dfrac{9~m^2/s^2}{1.5~m}=6~m/s^2\)

OpenStudy (anonymous):

Yes, I got the identical answer to you. Yet the options I'm given are: 3.0, 0.75, 1.0, and 9.0. None of which make any sense to me. @aaronq

OpenStudy (aaronq):

answers in the back of books often have mistakes :S but yeah it's good.

OpenStudy (anonymous):

Okay, thanks, I'll have to tell my teacher about it.